# Design and Verification of the JUNO Liquid Filling Control System

### v1

@@ Abstract
Jiangmen Underground Neutrino Observatory (JUNO) is a large-scale neutrino experiment with multiple physics goals including neutrino mass hierarchy, accurate measurement of neutrino oscillation parameters, neutrino detection from supernova, sun, and earth, etc. The Central Detector (CD) of JUNO using 20-kiloton Liquid Scintillator (LS) as target mass with 3% energy resolution at 1 MeV and low radioactive background. To achieve LS filling smoothly and detector running safely, JUNO’s liquid Filling, Overflowing, and Circulating System (FOC) features a control system based on a Programmable Logic Controller and equipped with high-reliable sensors and actuators. This paper describes the design of the FOC automatic monitoring and control system including hardware and software. The control logic including Proportional-Integral-Derivative control, sequential control, and safety interlocks ensures precise regulation of key parameters such as flow rate, liquid level, and pressure. Test results demonstrate the FOC system can satisfy the requirement of JUNO detector filling and running operations.
@@ 1 Introduction
The filling, overflowing, and circulating (FOC) system is in charge of LS filling, overflowing or refilling, and online circulation for LS re-purification [2]. Filling the CD with purified LS is the final step in the detector construction. The process must preserve the structural integrity of the detector and the chemical and optical properties of the LS. During detector filling and data taking, the system should also be able to maintain LS pressure and control the LS level for detector safety. To meet the above requirements, a FOC system is designed with an operation lifetime of 20 years.
@@
This paper focuses on the control system of the FOC system. The paper is organized as follows. In Section 2, we describe the design, specifications, and principles related to the FOC control system. In Section 3, we describe the principles of control logic and their application scenarios. We present the test performances of the control system in Section 4.
@@ 2.1 System requirement
The filling process consists of: (1) synchronous pure water injection into both the CD and WP, (2) replacing pure water in the CD with LS. During the data-taking period, the system must maintain stable CD pressure while limiting liquid level fluctuations to within 20 cm (overflow and refilling condition), and support online LS circulation. To ensure these functions, the FOC control system must meet several major requirements: precision, robustness, scalability, stability, anti-interference, safety and long time data traceability.
Firstly, the FOC control system must exhibit extremely high accuracy. The measurement deviations of all sensors shall be strictly constrained within the allowed range permitted by the process to ensure that every parameter during the filling and detector running accurately reflects the system’s status. High-precision sensors provide a solid foundation for precise system control, with specific accuracy requirements such as a liquid level sensor measurement accuracy $\leq 0.2\%$ full scale (FS) and a flow sensor accuracy reaching ±$\pm 0.1\%$ FS. The total control shall maintain a deviation $\leq 1\%$ between the setpoint and feedback value in a closed-loop control. Pumps shall maintain rotational speed stability within ±$\pm 0.2\%$ of setpoint with $\leq 10$ second standby activation. Regulating valves require ±$\pm 1\%$ positioning accuracy, while on-off valves must achieve < 1 second response time and $<10^{-6}\,\text{mbar}\cdot\text{L}\cdot\text{s}^{-1}$ leak rate.
Secondly, the FOC control system must possess security and scalability. Specifically, it requires safety interlock mechanisms to mitigate risks to detectors during abnormal conditions. A physical emergency stop function is needed to ensure process termination even during system failures. For safety, we also use a "dual-confirmation" protocol to prevent hazardous operations caused by mistakes. Considering potential changes and upgrades in FOC system, the control system must feature modular expandability to avoid system redesigns.
In the end, the FOC control system requires high stability, robustness and anti-interference capability. It must maintain precise control during prolonged continuous operation to avoid risks caused by control deviations and equipment failures. The system should be able to handle environmental fluctuations (e.g., electromagnetic interference, temperature variations, mechanical vibrations) or internal signal noise.
@@ 2.2 System design
Based on the above requirements, the layout of the FOC control system is shown in Figure 2. The system integrates 35 sensors, 20 control valves, and 8 pump units, forming a sophisticated control network. In this configuration, the system manages more than 80 sets of analog input/output (I/O) control quantities and more than 150 digital I/O control and alarm signals. Therefore, it mainly consists of 4 layers: the sensor layer, the controller layer, the actuator layer, as well as the alarm and data storage layer.
Figure 2: Layout of FOC control system
@@ 2.2.1 Sensor layer
As the critical monitoring terminal, all sensors selected for the FOC system (including level, pressure, temperature, flow rate, etc.) are sourced from Endress+Hauser (E+H) [7], renowned for their stability and reliability with an accuracy better than 0.2% FS. The instrument selection has undergone multiple rounds of demonstration, aligning with the system pipeline or tanks design and meeting the corresponding filling index requirements. For instance, the flowmeters were specifically chosen to match the LS filling rate of $7\ \text{m}^{3}/\text{h}$ with optimal measuring range and precision. These devices have undergone rigorous acceptance testing, including factory tests, offline tests, and online tests. All sensors come pre-equipped with functions such as preset alarm thresholds, alarm signal output, and fault diagnostics. The surface roughness of sensor interfaces in contact with LS meets design requirements and maintains sufficient cleanliness.
@@
Sensor type Quantity Accuracy (FS) Purpose
Electronic remote differential pressure level gauge 4 0.20% Measure CD liquid level (vertical tube & horizontal tube)
Laser level gauge 1 0.10% Measure CD liquid level
Differential pressure level gauge 2 0.20% Measure liquid level in the upper chimney
Static pressure liquid level gauge 5 0.20% Measure WP liquid level (bottom & surface)
Differential pressure level gauge 5 0.20% Measure tank liquid level
Radar wave level gauge 3 0.10% Measure tank liquid level
Temperature sensor 3 0.20% Measure LS temperature inside tanks
Coriolis mass flowmeter 2 0.10% Measure flow of exchanging LS/water
Pressure transmitter 3 0.15% Measure tank gas pressure
Pressure transmitter 3 0.15% Measure pump discharge pressure
Pressure transmitter 1 0.20% Measure pressure of calibration house
Double-flange pressure transmitter 1 0.10% Measure pressure difference before and after the filter
Vortex flowmeter 2 0.20% Measure water flow during water filling period
The sensor specifications are summarized in Table 1, which aligns with the system’s requirements. Specifically, redundant sensors are implemented at the critical points: (1) CD level monitoring uses 2 pairs electronic remote differential pressure gauges, 1 laser gauge, and 2 chimney differential pressure gauges; (2) WP employs 5 static pressure gauges at different heights; (3) tanks feature dual level sensors; (4) nitrogen pressure is measured by 3 sensors at 3 connected tanks; (5) four temperature sensors are distributed throughout LS pipelines.
@@ 2.2.2 Controller layer
The Siemens S7-300 series Programmable Logic Controller (PLC) is chosen as the core component of the FOC control system. Equipped with a high-performance CPU, it has powerful computing and processing capabilities. It can analyze and judge various input signals in micro-second and then quickly issue control commands including equipment start-stop operations and process parameter adjustments. Its modular design enables flexible configuration to meet diverse control requirements while ensuring system reliability through independent module replacement, maintaining uninterrupted operation.
@@
Control software was developed based on Totally Integrated Automation (TIA) Portal [8], which integrates digital planning, operation management, diagnostic capabilities and comprehensive monitoring of the FOC system. The system can acquire and display various parameters in real-time. In addition, we can operate various devices through the control interface, such as adjusting valve openings and controlling pump frequency.
For real-time interaction among different control systems in JUNO, another set of monitoring and control programs is developed based on the Experimental Physics and Industrial Control System (EPICS) [9] framework, which provide functions such as comprehensive monitoring, historical query, and data caching. Moreover, the heartbeat packet program is used to monitor the operational status of PLC. The human-machine interface (HMI) developed based on TIA Portal is shown in Figure 3.
@@ 2.2.3 Actuator layer
All electromagnetic pumps operate in a “one operation and one standby” configuration. Real-time pump status monitoring is achieved through digital feedback signals (ready, fault, running) and analog feedback of actual operating frequency, which are compared against control signals. Pressure sensors downstream of pumps monitor hydraulic pressure to prevent overpressure, while upstream level interlocks safeguard against dry running. All valves come with ample spare parts and manual bypass valves. On-off valves provide “opened” or “closed” status feedback signals, while regulating valves offer analog feedback on opening degrees. Both pumps and valves have additional monitoring methods include local inverter displays and regular operator rounds, ensuring system stability.
@@ 2.2.4 Alarm and data management layer
In terms of data management, all data related to FOC are properly stored. The system employs a dual data storage architecture, with local hard disk storage enabling real-time analysis and troubleshooting, while synchronized Detector Control System (DCS) storage facilitates distributed access and multi-department data sharing. This integrated approach, coupled with web-based remote monitoring capabilities, ensures comprehensive data traceability.
@@ 3.1 Architecture of the control logic
Building on the hardware architecture described in Section 2.2, drawing on the hierarchical control theory ISA-88 [13], the FOC control system employs a closed-loop control framework through three functional layers: detection, control, and monitoring & safety. This hierarchical design enables precise regulation of LS parameters while maintaining system stability during long-term operation.
@@
    This layer interfaces with high-reliability sensors to collect real-time process parameters, employing a hybrid filtering strategy to mitigate noise. Arithmetic mean filtering reduces high-frequency fluctuation such as radar level gauge noise from 30 mm to 5 mm, while time hysteresis filter combines dual conditions of threshold and duration, avoiding false triggers caused by transient interference signals of Coriolis flow meters.
@@ 3.2 Typical control logic types and application scenarios
    Taking LS filling control as an example: the PID controller dynamically adjusts the flow through three coordinated actions, when actual flow falls below the setpoint, the proportional term immediately boosts output. The derivative term counters rapid flow increases by generating opposing signals based on the rate of deviation change, preventing overshoot. Simultaneously, the integral term continuously eliminates residual errors - during initial filling, strong proportional action dominates, while the integral term handles fine adjustments as flow nears the target, with derivative action maintaining stability throughout the transition.
@@
    The safety interlock logic mitigates accidents by sensing process parameters out of range and acting immediately. Risks during detector filling or operation, such as detector pressure exceeding the setpoint, abnormal LS exchanging flux or LS temperature, may damage the detector. The safety interlock logic following two basic principles: independence and fail-safety. For the independence principle, the safety logic operates independently of the normal control logic to avoid common-cause failures. For example, the detector calibration house pressure monitoring uses independent sensors in the safety interlock system, ensuring accurate monitoring and triggering of safety actions even if normal control sensors fail. The fail-safety principle means the system will start the safe action by default in case of failure. Critical equipment such as pumps and valves shut down automatically in case of CD pressure imbalance.
@@ 4.1 Test with pure water system
The test results (Fig. 4) demonstrate the rapid response and precise flow control of the FOC control system, with the WP ($39.3\ \text{m}^{3}/\text{h}$) and CD ($18.9\ \text{m}^{3}/\text{h}$) pipelines reaching the target flows in 2 minutes and maintaining stability within 0.5% uncertainty, meeting all design specifications.
@@ 4.2 Test with LS purification system
During test, LS was transferred from the purification system to FOC tanks at varying flow rates, with automatic pump control triggered by tank level thresholds. Figure 5 shows the changes in liquid level at $4.5\ \text{m}^{3}/\text{h}$ during the transfer of LS between tanks. Although the comparison between differential pressure and radar level gauges confirms measurement accuracy (deviation < 0.2%), some limitations exist: (1) the differential pressure gauge exhibited a 0-0.3 m dead zones due to installation constraints; (2) The radar gauge demonstrates both a 0.2 m dead zone (for FOC horizontal tanks) and measurable response lag.
@@ 4.3 Internal test of FOC system
Internal test has proved the robust, safe, and accurate performance of the control system in overflowing and LS exchanging scenarios.

### v3

@@ Abstract
Jiangmen Underground Neutrino Observatory (JUNO) is a large-scale neutrino experiment with multiple physics goals including neutrino mass hierarchy, accurate measurement of neutrino oscillation parameters, neutrino detection from supernova, sun, and earth, etc. This paper presents the design, implementation, and verification of a high-reliability automated control system for the liquid Filling, Overflow, and Circulation system in the JUNO experiment. The system is built upon a Programmable Logic Controller architecture, integrated with high-precision sensors and actuators. It implements advanced control strategies, including Proportional-Integral-Derivative regulation, sequential logic, and safety interlocks, to achieve closed-loop control of critical parameters such as flow rate, liquid level, and pressure. Commissioning tests with both pure water and liquid scintillator demonstrate the system’s exceptional performance, achieving flow control stability within 0.5% of the setpoint with a rapid stabilization time. The robust design, featuring hardware redundancy and software safeguards, ensures the system meets the stringent requirements for the safe filling and long-term stable operation of JUNO’s 20-kiloton central detector and provides a scalable reference for large underground fluid control experiments.
@@ 1 Introduction
The filling, overflow, and circulation (FOC) system is in responsible for LS filling, overflowing or refilling, and online circulation for LS re-purification [2]. Filling the CD with purified LS is the final step in the detector construction. The process must preserve the structural integrity of the detector and the chemical and optical properties of the LS. During detector filling and data taking, the system should also be able to maintain LS pressure and control the LS level for detector safety. To meet the above requirements, the FOC system is designed with an operation lifetime of 20 years.
@@
This paper focuses on the control system of the FOC system. The paper is organized as follows. In Section 2, we describe the design, specifications, and principles related to the control system. In Section 3, we describe the principles of control logic and their application scenarios. We present the test performances of the control system in Section 4.
@@ 2.1 System requirement
The filling process consists of: (1) synchronous pure water injection into both the CD and WP, (2) replacing pure water in the CD with LS. During the data-taking period, the system must maintain stable CD pressure while limiting liquid level fluctuations to within 20 cm (overflow and refilling condition), and support online LS circulation. To ensure these functions, the control system must meet several major requirements: precision, robustness, scalability, stability, anti-interference, safety and long time data traceability.
Firstly, the control system of FOC must exhibit extremely high accuracy. The measurement deviations of all sensors shall be strictly constrained within the allowed range permitted by the process to ensure that every parameter during the filling and detector running accurately reflects the system’s status. High-precision sensors provide a solid foundation for precise system control, with specific accuracy requirements such as the liquid level sensor measurement accuracy $\leq 0.2\%$ full scale (FS) and the flow sensor accuracy reaching $\pm 0.1\%$ FS. The system is designed to maintain deviations within $1\%$ between the setpoint and feedback value in a closed-loop control. Pumps shall maintain rotational speed stability within $\pm 0.2\%$ of setpoint with $\leq 10$ seconds standby activation. Regulating valves require $\pm 1\%$ positioning accuracy, while on-off valves must achieve < 1 second response time and $10^{-6}\,\text{mbar}\cdot\text{L}\cdot\text{s}^{-1}$ leak rate.
Secondly, the control system must possess security and scalability. Specifically, it requires safety interlock mechanisms to mitigate risks to detectors during abnormal conditions. A physical emergency stop function is needed to ensure process termination even during system failures. For safety, we also use a "dual-confirmation" protocol to prevent hazardous operations caused by mistakes. Considering potential changes and upgrades in FOC system, the control system must feature modular expandability to avoid system redesigns.
In the end, the control system requires high stability, robustness and anti-interference capability. It must maintain precise control during prolonged continuous operation to avoid risks caused by control deviations and equipment failures. The system should be able to handle environmental fluctuations (e.g., electromagnetic interference, temperature variations, mechanical vibrations) or internal signal noise.
@@ 2.2 System design
Based on the above requirements, the layout of the control system is shown in Figure 2. The system integrates 35 sensors, 20 control valves, and 8 pump units, forming a sophisticated control network. In this configuration, the system manages more than 80 sets of analog input/output (I/O) control quantities and more than 150 digital I/O control and alarm signals. Therefore, it mainly consists of 4 layers: the sensor layer, the controller layer, the actuator layer, as well as the alarm and data storage layer.
Figure 2: Layout of the control system
@@ 2.2.1 Sensor layer
[…]
@@
Sensor type Quantity Accuracy (%FS) Purpose
Electronic remote differential pressure level gauge 4 0.20 Measure CD liquid level (vertical tube & horizontal tube)
Laser level gauge 1 0.10 Measure CD liquid level
Differential pressure level gauge 2 0.20 Measure liquid level in the upper chimney
Static pressure liquid level gauge 5 0.20 Measure WP liquid level (bottom & surface)
Differential pressure level gauge 5 0.20 Measure tank liquid level
Radar wave level gauge 3 0.10 Measure tank liquid level
Temperature sensor 3 0.20 Measure LS temperature inside tanks
Coriolis mass flowmeter 2 0.10 Measure flow of exchanging LS/water
Pressure transmitter 3 0.15 Measure tank gas pressure
Pressure transmitter 3 0.15 Measure pump discharge pressure
Pressure transmitter 1 0.20 Measure pressure of calibration house
Double-flange pressure transmitter 1 0.10 Measure pressure difference before and after the filter
Vortex flowmeter 2 0.20 Measure water flow during water filling period
The sensor specifications are summarized in Table 1, which aligns with the system’s requirements. Specifically, redundant sensors are deployed at critical points as follows: (1) For CD level monitoring: 2 pairs of electronic remote differential pressure gauges, 1 laser gauge, and 2 chimney differential pressure gauges; (2) For WP level monitoring: 5 static pressure gauges installed at different heights; (3) For tank level monitoring: dual level sensors for each tank; (4) For nitrogen pressure monitoring: 3 sensors installed at 3 connected tanks; (5) For LS temperature monitoring: four temperature sensors distributed throughout the pipelines and tanks.
@@
Besides, a rigorous multi-stage calibration and acceptance procedure was implemented for all sensors. First, each device was supplied with factory calibration certificates, and their documented accuracy and stability met the experiment’s stringent requirements. Then, our acceptance process comprised two phases: an off-line acceptance without load, and an on-line acceptance after installation. During these tests, the calibration of all sensors was verified. For those that could be calibrated within the FOC system itself—such as tank level gauges, pressure transmitters, temperature sensors, and pump discharge pressure sensors—we employed cross-comparison methods. A typical example is using two independent level gauges to validate each other’s readings. The success of these calibration checks is further demonstrated by the system’s performance during integrated testing, as detailed in Section 4. For critical sensors that could not be fully calibrated on-site prior to the JUNO detector filling—specifically, the CD and WP level gauges—a qualitative validation was performed using a small-scale mock-up. This setup allowed for cross-checking among multiple sensor groups, verifying their consistent response and functionality, thereby providing confidence in their readiness for the final operation.
@@ 2.2.2 Controller layer
The Siemens S7-300 series Programmable Logic Controller (PLC) is chosen as the core component of the control system. Equipped with a high-performance CPU, it has powerful computing and processing capabilities. It can analyze and judge various input signals in micro-second and then quickly issue control commands including equipment start-stop operations and process parameter adjustments. Its modular design enables flexible configuration to meet diverse control requirements while ensuring system reliability through independent module replacement, maintaining uninterrupted operation.
@@
The control software was developed using the Siemens Totally Integrated Automation (TIA) Portal [8], which provides integrated planning, diagnostic, and comprehensive monitoring capabilities. The primary human-machine interface (HMI) for local operation and real-time control is also developed and hosted within the TIA Portal environment, running on a dedicated industrial computer. This TIA-based HMI (shown in Figure 3) is the main interface for operators to perform control actions, such as adjusting valve openings and controlling pump frequency, and for displaying real-time process parameters during routine operations.
@@
For integration into the broader JUNO Detector Control System (DCS), a separate monitoring and control program is developed based on the Experimental Physics and Industrial Control System (EPICS) [9] framework. The PLC, the TIA Portal HMI station, and the EPICS Input/Output Controller (IOC) – running on a Linux-based industrial computer – are interconnected via a dedicated industrial Ethernet switch.
The data synchronization between the PLC and the EPICS IOC is established through a network communication protocol (e.g., OPC UA or native Siemens S7), allowing the EPICS system to access all critical process variables. While the TIA HMI remains the primary interface for direct control, the EPICS channel serves two pivotal functions:
  • It enables comprehensive monitoring, historical data querying, and alarm management for the FOC system from the unified JUNO-wide DCS interface, facilitating coordinated operation with other subsystems.
  • It implements a heartbeat monitoring program. The EPICS IOC continuously checks the operational status of the PLC. A loss of this heartbeat signal triggers a global alarm within the DCS, notifying operators of a potential PLC failure.
This dual-interface architecture ensures both robust, low-latency local control via the TIA ecosystem and seamless integration into the experiment’s global monitoring and safety framework via EPICS.
@@ 2.2.3 Actuator layer
All electromagnetic pumps operate in the "one operation and one standby" configuration. Real-time pump status monitoring is achieved through digital feedback signals (ready, fault, running) and analog feedback of the actual operating frequency, which are compared against control signals. Pressure sensors downstream of pumps monitor hydraulic pressure to prevent overpressure, while upstream level interlocks safeguard against dry running. All valves come with ample spare parts and manual bypass valves. On-off valves provide “opened” or “closed” status feedback signals, while regulating valves offer analog feedback on opening degrees. Both pumps and valves have additional monitoring methods include local inverter displays and regular operator rounds, ensuring system stability.
The actuators underwent rigorous acceptance testing to verify their performance aligns with the control system’s requirements. Regulating valves were characterized to establish their flow coefficient curves, ensuring the PID control loops can accurately relate valve opening percentage to flow rate. The positioning feedback of these valves is continuously monitored, and any significant deviation from the commanded signal triggers a maintenance alert.
Similarly, the electromagnetic pumps were tested to confirm their rotational speed stability and response time under various load conditions. This characterization process for both valves and pumps is intrinsically coupled with the system performance tests described in Section 4. The stable flow control and precise pressure maintenance achieved during those tests (e.g., as summarized in Table 3) serve as the ultimate validation of the actuator calibration and integration. This proactive approach, combined with real-time status monitoring and the "one operation and one standby" configuration, minimizes the uncertainty in system actuation and provides a high-degree of operational reliability.
@@ 2.2.4 Alarm and data management layer
In terms of data management, all data related to FOC are properly stored. The system employs a dual data storage architecture, with local hard disk storage enabling real-time analysis and troubleshooting, while synchronized DCS storage facilitates distributed access and multi-department data sharing. This integrated approach, coupled with web-based remote monitoring capabilities, ensures comprehensive data traceability.
@@ 3.1 Architecture of the control logic
Building on the hardware architecture described in Section 2.2, drawing on the hierarchical control theory ISA-88 [13], the control system employs a closed-loop control framework through three functional layers: detection, control, and monitoring & safety. This hierarchical design enables precise regulation of LS parameters while maintaining system stability during long-term operation.
@@
    This layer interfaces with high-reliability sensors to collect real-time process parameters, employing a hybrid filtering strategy to mitigate noise. Arithmetic mean filtering reduces high-frequency fluctuations such as radar level gauge noise from 30 mm to 5 mm, while the time hysteresis filter combines dual conditions of threshold and duration, avoiding false triggers caused by transient interference signals of Coriolis flow meters.
@@ 3.2 Typical control logic types and application scenarios
    Taking LS filling control as an example: the PID controller dynamically adjusts the flow through three coordinated actions. When the actual flow falls below the setpoint, the proportional term immediately boosts the output. The derivative term counters rapid flow increases by generating opposing signals based on the rate of deviation change, preventing overshoot. Simultaneously, the integral term continuously eliminates residual errors - during initial filling, strong proportional action dominates, while the integral term handles fine adjustments as flow nears the target, with derivative action maintaining stability throughout the transition.
@@
    The safety interlock logic mitigates accidents by sensing process parameters out of range and acting immediately. Risks during detector filling or operation–such as detector pressure exceeding the setpoint, abnormal LS exchange flux, or abnormal LS temperature–may damage the detector. The safety interlock logic follows two basic principles: independence and fail-safety, which are aligned with the guidance of functional safety standards such as IEC 61508 [15] and IEC 61511 [16].
    For the independence principle, the safety logic operates independently of the normal control logic to avoid common-cause failures. For example, the detector calibration house pressure monitoring uses independent sensors in the safety interlock system, ensuring accurate monitoring and triggering of safety actions even if normal control sensors fail. The fail-safety principle means the system will start the safe action by default in case of failure. Critical equipment such as pumps and valves shut down automatically in case of CD pressure imbalance.
    Furthermore, the selection of specific equipment also incorporates inherent safety interlock features, thereby reinforcing the overall system’s reliability. A prime example is the use of ABB ACS580 [17] variable frequency drives for the electromagnetic pumps. These drives provide a critical, equipment-level layer of protection by continuously monitoring motor current and speed. Should any parameter exceed its predefined safe operating limit, the drive will immediately initiate a safe stop, independently of the higher-level control system. This built-in functionality embodies the fail-safety principle directly at the actuator level and complements the system-wide interlock architecture.
@@ 4.1 Test with pure water system
[…]
@@ 4.2 Test with LS purification system
[…]
@@ 4.3 Internal test of FOC system
The performance of the control system across all tested scenarios is quantitatively summarized in Table 3. The results demonstrate the system’s capability to meet or exceed the design specifications outlined in Section 2.1. Key achievements include the rapid stabilization of flow rates within 1–3 minutes and the maintenance of control accuracy within 0.5% under both water and LS conditions. The LS transfer test further verified the high precision of the sensor suite, with level gauge deviations remaining below 0.2% of the full scale. These comprehensive tests confirm the system’s robustness, safety, and precision, establishing its readiness for the JUNO detector filling and long-term operation.
Table 3: Summary of key performance metrics from tests
Test Scenario Parameter Performance Summary
Pure Water Filling (WP) Flow Rate
Setpoint: $39.3~\mathrm{m^{3}/h}$, Measured: $39.3\pm 0.2~\mathrm{m^{3}/h}$
Deviation: $\leq 0.5\%$, Stabilized in $1~\mathrm{min}$
Pure Water Filling (CD) Flow Rate
Setpoint: $18.9~\mathrm{m^{3}/h}$, Measured: $18.9\pm 0.2~\mathrm{m^{3}/h}$
Deviation: $\leq 0.5\%$, Stabilized in $1~\mathrm{min}$
LS Transfer Liquid Level
Setpoint: $1500~\mathrm{mm}$, Measured: $1500\pm 3~\mathrm{mm}$
Deviation: $0.2\%$ FS
LS Filling Flow Rate
Setpoint: $1500~\mathrm{L/h}$, Measured: $1500\pm 10~\mathrm{L/h}$
Deviation: $\leq 0.5\%$, Stabilized in $3~\mathrm{min}$